OpenStudy (anonymous):

If the area of a circle is 8pi , what is the length of its diameter?

OpenStudy (angela210793):

A=pir^2 8pi=pir^2 r^2=8 r=2sqr2 d=2r=2*2sqr2=4sqr2

OpenStudy (lalaly):

area=pi r^2 8pi = pi r^2 8=r^2 r=sqrt8 diameter = r*2 = 2*sqrt8 = 4sqrt2

OpenStudy (anonymous):

5.65

OpenStudy (anonymous):

yup thats the one!!!...haha thanks....im gonna post some more questions :)

OpenStudy (anonymous):

thanks angela and lalay

OpenStudy (angela210793):

yw :)